OpenStudy (anonymous):

Square root of 48x^3 divided by the square root of 3xy^2

OpenStudy (xishem):

\[\frac{\sqrt{48x^3}}{\sqrt{3xy^2}}=\frac{4x \sqrt{3x}}{y \sqrt{3x}}=\frac{4x}{y}\]

OpenStudy (saifoo.khan):

Rite^

OpenStudy (anonymous):

cant i do 48/3=\[\sqrt{16x^{2}}y ^{2}= 4xy\]

OpenStudy (xishem):

\[\frac{48}{3}\neq \frac{\sqrt{48}}{\sqrt{3}}\]

OpenStudy (xishem):

\[16 \neq 4\]